The Language is Java

Directions:

  1. Answer the question for each problem,
  2. Trace the problem as part of the answer.

Arrays and ArrayLists #2

public void a2(int[] a, int n)

{

   if (n > 0)

   {

      a[n - 1] += a[n];

      a2(a, n - 1);

   }

}

int[] arr = {1, 2, 3, 4};

What is arr after a2(arr, arr.length - 1) ?
Answer #1
int[] arr = {1, 2, 3, 4};
What is arr after a2(arr, arr.length - 1) ?

arr[2] = arr[2] + arr[3] = 3 + 4 = 7
arr[1] = arr[1] + arr[2] = 2 + 7 = 9
arr[0] = arr[0] + arr[1] = 1 + 9 = 10

So, the final content of arr is {10, 9, 7, 4}